OpenStudy (anonymous):

Solve the following: One car left Carville driving toward Beattyville, which is 300 miles away. A second car left Beattyville at the same time, driving toward Carville. If the first and second cars are averaging 55 mph and 65 mph, respectively, in how many hours will they pass each other? 2 hours 3.5 hours 2.5 hours 3 hours

OpenStudy (anonymous):

lets check an answer and see if it works. suppose \(t=2\) so they each travel for 2 hours. since distance = rate times time, one car would have gone \(2\times 55=110\) miles, the other would have gone \(2\times 65=130\) miles. so the total would be \(110+130=240\) miles, and so that is not correct since the total has to be 300

OpenStudy (anonymous):

this suggests a way to solve the problem one car travels \(55\times t\) miles and the other travels \(65\times t\) miles and the total distance traveled must be 300 so we can set \[55t+65t=300\] and solve for t

OpenStudy (anonymous):

can you solve that one?

OpenStudy (anonymous):

10